OpenStudy (nightmarenight):

Help??? Will fan and give medal for correct answer! Find k(4) for the function k(n) = 3n^2.

OpenStudy (anonymous):

\[k(n) = 3n^2\] \[k(4) = 3(4)^2\] \[k(4) = 3*16=48\]